OpenStudy (anonymous):

HELP PLEASE ! ! ! ! ! ! ! ! ! 2x-y+2z=-7 -x+2y-4z=5 x+4y-6z=-1 Solve The System Using The Elimination Method.

OpenStudy (anonymous):

The solution is x = -3, y = -1, z = -1

OpenStudy (anonymous):

Could You Please Explain How That Worked ?

OpenStudy (anonymous):

* How You Got That Answer

jhonyy9 (jhonyy9):

assume second with 3rd and you get 6y-10z=4

OpenStudy (anonymous):

What About The Second Equation ? I Got 6y-10z=4 & 3y-6z=3 As My Two Equations .

jhonyy9 (jhonyy9):

multiply second equation with 2 and assum it with first equation so -2x+4y-8z=10

jhonyy9 (jhonyy9):

2x-y+2z=-7 -2x+4y-8z=10 -------------- 0 3y-6z=3 / divide by 3 y-2z=1 y-2z=1 6y-10z=4 /divide by 2 and will get 3y-5z=2 y-2z=1 multiply by 3 and subtract second from first equation 3y-5z=2 3y-6z=3 3y-5z=2 --------- 0 -z=1 z=-1 y-2z=1 y-2(-1)=1 y+2=1 y=-1 x=2y-4z-5 x=2(-1)-4(-1)-5=-2+4-5=-7+4=-3 x=-3 y=-1 z=-1